OpenStudy (anonymous):

what is the composite function (SoP)(y) where S(x)=x^2 & P(x)=2^x the answer should be in number

OpenStudy (anonymous):

my answer comes 2^2x :(

OpenStudy (anonymous):

sorry 2^2y

OpenStudy (anonymous):

2^2y is correct.

OpenStudy (anonymous):

then why in text this condition is mentioned "In each case your answer should be a number"?

OpenStudy (anonymous):

similarly for S(x)=x^2 and s(x)=sinx then (Sos)(y)=sin^2 y

OpenStudy (anonymous):

this is not a number but the answer is written in form of sin^2 y in the solution mannual

OpenStudy (anonymous):

Maybe you can learn something more about y?

OpenStudy (anonymous):

y is not given

OpenStudy (anonymous):

it's in Spivak

OpenStudy (anonymous):

What's Spivak?

OpenStudy (anonymous):

Spivak's calculus book which is more of analysis

OpenStudy (anonymous):

Ok, in any case, you can't do any better than 2^2y